OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 5x + 35y = -10 15x + 5y = 170

OpenStudy (anonymous):

5x = -35y - 10 x = -7y - 2 15x = -5y +170 x = -1/3y + 34/3 -7y -2 = -1/3y + 34/3 -2 = 6.67y + 34/3 -13.33 = 6.67y y = -2

OpenStudy (anonymous):

5x = -35 (-2) -10 5x = 70 -10 5x = 60 x = 12